# Organic or Diffused: Can We Distinguish Human Art from AI-generated Images?

### v2

@@ 6.1. Study Setup
Participants.    As discussed in §3.3, we recruited three participant groups, and a total of 3993 participants completed our study and passed all attention/consistency checks. These include 177 baseline participants recruited from Prolific, referred to as general users, 3803 professional artist volunteers, and 13 high profile experts.
@@
After a brief introduction on generative models and the current issue of distinguishing between human artworks and AI-generated images, we present the participant with a sequence of 20 images, shown one at a time, and ask them to decide whether each image is human-made or AI-generated. For each image, we ask two questions. The first question is “Have you seen this image before?” If yes, we disregard the response for this image, since the participant has likely seen this image and is possibly aware of the image’s true source. The second question asks them to rate the current image with one of the five choices: “human art (very confident),” “human art (somewhat confident),” “not sure,” “AI-generated (somewhat confident),” “AI-generated (very confident).” For each participant, we randomly sample 20 images from a collection of 670 images, consisting of 210 human artworks, 70 human artworks after upscaling, 350 AI-generated images, and 40 hybrid images.
Next, we ask the participant, for each of the 7 art styles, whether they are confident at distinguishing human art from AI-generated images. If confident, we ask them to describe the properties of the art that contributed to their decisions. Here we present seven options: “content,” “complexity,” “technical skill,” “perspective,” “lighting,” “consistency,” and an additional text response section for additional details. Finally, we ask them if they self-identify as full-time or part time artists and why.
@@ 6.4. Detection Accuracy for Expert Artists
Our 13 expert artists show greater proficiency in the detection task, raising overall detection accuracy (ACC) to 83% (see last row in Table 5). They also produce slightly imbalanced FPR (20.78%) and FNR (14.63%), indicating that they are better at spotting AI-generated images than human art. When we gave feedback to the experts on their results, many were frustrated that they committed errors by marking human art as AI (false positive). In retrospect, they explained that they identified detailed mistakes and inconsistencies which were likely due to inexperience and human error by the human artists. For example, in a painting of a bedroom bathed in moonlight, the shadow of a window pane had slightly offset position of the latch compared to the window itself. This was seen by expert artists as an inconsistency, but later attributed to lack of attention to detail by a human artist.
@@
Focus Study on “Hard-to-Detect” AI Images.    For a comprehensive view of how experts identify AI-generated images, we presented our expert group with a fixed set of the six AI-generated images that produced the most false negative errors by the professional (non-expert) artist group. They cover 5 styles: sketch, oil/acrylic, cartoon, watercolor, and anime. For each of these difficult images, we asked the experts for detailed feedback on what exactly they would use to identify these images as AI. Figure 4 shows all six images, each followed by two regions of the image zoomed in to show details of the artifacts identified by our expert artists.
@@ 8. Combining Human and Automated Detectors
Our study shows that both human artists and automated detectors face challenges in distinguishing between human art and AI-generated images. Tools like Hive are highly effective at evaluating unperturbed images, but perform poorly when AI-generated images are intentionally perturbed (e.g., by applying Glaze) to evade detection. On the other hand, human experts can still identify AI-generated images when they are Glazed. Thus we believe a more effective solution is to form a mixed team of human artists and machine classifiers.
Teaming up Hive and Expert Artist.    We evaluate the scenario of teaming up Hive with one expert artist for each image. Now the combined team has two “judges”. In case of a tie, we use the decision with higher confidence score, i.e. picking a “very confident” decision over a “somewhat confident” one.
@@ 9. Ethics
Our user study protocol was reviewed and approved by our institutional review board (IRB). In our study design, we prioritized consent and protection of all participants, particularly human artists and their artworks.
@@ 10. Discussion and Takeaways
Distinguishing human art from the output of generative AI models is an important problem whose impact will only grow with time. This paper describes the results of our study to determine the efficacy of today’s tools to address this challenge, including detection via automated tools (supervised classifiers), detection systems from research, and detection by human artists.
Our results include several counter-intuitive findings with longer term implications. We find that supervised classifiers can be quite accurate, but are weaker on input for which it has limited training data on, e.g. newer models and adversarially altered images. These limitations are likely to persist with arrival of new generators and adversarial perturbation methods. In contrast, human artists (especially experienced experts) do well across all inputs. While experts still fall short of Hive in accuracy, they are more robust to perturbed images, and their errors often arise from identifying human artist mistakes as errors made by AI models. Our results show that a mixed team of human and automated detectors will likely provide the best combination of detection accuracy and robustness to adversarial measures.
@@
Additional Information on Data Collection.    Table 12 summarizes the number of images curated as human artworks for each of the 7 art styles and the AI-generated images per generation model. Table 13 lists the modifications made to the extracted BLIP captions, which are then used to prompt individual AI generators. And Figure 7 shows visual examples of the five types of perturbations considered in our study.
@@
Real Artworks AI-Generated Images
Human Artists CIVITAI DALL-E 3 Firefly MJv6 SDXL
Anime 40 10 10 10 10 10
Cartoon 40 10 10 10 10 10
Fantasy 40 10 10 10 10 10
Oil/Acrylic 40 10 10 10 10 10
Photography 40 10 10 10 10 10
Sketch 40 10 10 10 10 10
Watercolor 40 10 10 10 10 10
Total 280 70 70 70 70 70
350
Table 12. Our dataset includes 280 real artworks and 350 AI-generated images, covering 7 art styles and 5 AI generators. Each numerical value represents the # of images.

### v3

@@ 6.1. Study Setup
Participants.    As discussed in §3.3, we recruited three participant groups, and a total of 3993 participants completed our study and passed all attention checks. These include 177 baseline participants recruited from Prolific, referred to as general users, 3803 professional artist volunteers, and 13 high profile experts.
@@
After a brief introduction on generative models and the current issue of distinguishing between human artworks and AI-generated images, we present the participant with a sequence of 20 images, shown one at a time, and ask them to decide whether each image is human-made or AI-generated. For each image, we ask two questions. The first question is “Have you seen this image before?” If yes, we disregard the response for this image, since the participant has likely seen this image and is possibly aware of the image’s true source. The second question asks them to rate the current image with one of the five choices: “human art (very confident),” “human art (somewhat confident),” “not sure,” “AI-generated (somewhat confident),” “AI-generated (very confident).” For each participant, we randomly sample 20 images from a collection of 670 images, consisting of 210 human artworks, 70 human artworks after upscaling, 350 AI-generated images, and 40 hybrid images. Every image is seen by five participants.
Next, we ask the participant, for each of the 7 art styles, whether they are confident at distinguishing human art from AI-generated images; the user study interface is presented in Figure 6. If confident, we ask them to describe the properties of the art that contributed to their decisions. Here we present seven options: “content,” “complexity,” “technical skill,” “perspective,” “lighting,” “consistency,” and an additional text response section for additional details. Finally, we ask them if they self-identify as full-time or part time artists.
@@ 6.4. Detection Accuracy for Expert Artists
Our 13 expert artists show greater proficiency in the detection task, raising overall detection accuracy (ACC) to 83% (see last row in Table 5). They also produce slightly imbalanced FPR (20.78%) and FNR (14.63%), indicating that they are better at spotting AI-generated images than human art. When we gave feedback to the experts on their results, many were frustrated that they committed errors by marking human art as AI (false positive). In retrospect, they explained that they identified detailed mistakes and inconsistencies which were likely due to inexperience and human error by the human artists. For example, in a painting of a bedroom bathed in moonlight, the shadow of a window pane had slightly offset position of the latch compared to the window itself. This was seen by expert artists as an inconsistency, but later attributed to lack of attention to detail by a human artist. This attests for the drop in ADSR for Firefly images. Since expert artists look at fine-grained detail in AI-generated images, they are overfitted to spot irregularities from popular generators and have yet been accustomed to the newer style of images.
@@
Focus Study on “Hard-to-Detect” AI Images.    For a comprehensive view of how experts identify AI-generated images, we presented our expert group with a fixed set of the six AI-generated images that produced the most false negative errors by the professional (non-expert) artist group. They cover 5 styles: sketch, oil/acrylic, cartoon, watercolor, and anime. For each of these difficult images, we asked the experts for detailed feedback on what exactly they would use to identify these images as AI. Figure 7 shows all six images, each followed by two regions of the image zoomed in to show details of artifacts identified by our expert artists.
@@ 8. Combining Human and Automated Detectors
Our study shows that both human artists and automated detectors face challenges in distinguishing between human art and AI-generated images. Tools like Hive are highly effective at evaluating unperturbed images, but perform poorly when AI-generated images are intentionally perturbed (e.g., Glaze or image overlays) to evade detection. On the other hand, human experts can still identify perturbed AI-generated images. Thus we believe a mixed team of human artists and machine classifiers will be the most effective.
Teaming up Hive and Expert Artists.    We evaluate a scenario that combines Hive scores with one expert artist. If Hive and the expert disagree, the score with higher confidence wins.
@@ 9. Ethics
Our user study was reviewed and approved by our institutional review board (IRB). In our study, we prioritized consent and protection of all participants, especially human artists and their artworks.
@@ 10. Discussion and Takeaways
As with any real-world study, there are limitations in our study that need to be considered.
  • Category of styles: our dataset only included a few fixed art styles. More diverse styles might provide more comprehensive results.
  • Cropping of images: cropping was applied to a small number of human artworks. We avoided samples with highly irregular aspect ratios and ensured meticulous cropping to minimize irregularities.
  • Curating Likert Scale: we discarded “not sure” responses from the user study. This was done to maintain consistency with measuring the confidence level (40-60%) in automated detectors and to prevent guesses from affecting our metrics.
Our results also suggest takeaways for different audiences.
For artists.    Proving human authenticity will be come increasingly important, and increasingly difficult. No single method wil be foolproof, and artists should consider incorporating work-in-progress (WIP) or timelapses into their process.
For researchers.    Developing highly accurate detection tools requires continued investment in ethically obtaining diverse training sets. To enhance robustness, incorporating perturbed and adversarially altered images during training is crucial.
For policy makers.    Implementing standards that mandate mixed detection teams in critical applications can enhance both detection accuracy and robustness.
As AI evolves, generative AI users seeking to avoid detection will adapt to exploit vulnerabilities. Some are using Adobe’s Photoshop to add imperfections to avoid the smooth AI finishing look. Others are using one model to generate foreground objects and another to create backgrounds (Deighton, 2024). These evolving techniques will continue to present challenges for future AI detection systems.
Acknowledgements
We thank our anonymous reviewers for their insightful feedback. Sincere thanks also go to the thousands of artists who participated in our user study. This work is supported in part by NSF grants CNS-2241303 and CNS-1949650. Opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of any funding agencies.
@@ Acknowledgements
Additional Information on Data Collection.    Table 12 lists the modifications made to the extracted BLIP captions, which are then used to prompt individual AI generators. Figure 10 shows examples of five types of perturbations considered in our study.